A pharmacy has determined that a healthy person should receive 70 units of proteins, 100 units of carbohydrates and 20 units of fat daily. If the store carries the six types of health food with their ingredients as shown in the table below, what blend of foods satisfies the requirements at minimum cost to the pharmacy? Make the mathematical mod


1
Expert's answer
2022-02-02T09:07:44-0500

Question:

A pharmacy has determined that a healthy person should receive 70 units of proteins, 100 units of carbohydrates and 20 units of fat daily. If the store carries the six types of health food with their ingredients as shown in the table below, what blend of foods satisfies the requirements at minimum cost to the pharmacy? Make a mathematical model for the given problem

Foods Protein units Carbohydrates units Fat units Cost per unit

A 20 50 4 2

B 30 30 9 3

C 40 20 11 5

D 40 25 10 6

E 45 50 9 8

F 30 20 10 8


Solution:

minimize cost:

"z=2x_1+3x_2+5x_3+6x_4+8x_5+8x_6"

subject to:

"20x_1+30x_2+40x_3+40x_4+45x_5+30x_6\ge70" : amount of protein

"50x_1+30x_2+20x_3+25x_4+50x_5+20x_6\ge100" : amount of carbohydrate

"4x_1+9x_2+11x_3+10x_4+9x_5+10x_6\ge20" : amount of fat

where x1, x2, x3, x4, x5, xare units of 6 foods


solution using Simplex method:

After introducing artificial variables:

Min Z=2x1+3x2+5x3+6x4+8x5+8x6+0S1+0S2+0S3+MA1+MA2+MA3

subject to

20x1+30x2+40x3+40x4+45x5+30x6-S1+A1"\ge" 70

50x1+30x2+20x3+25x4+50x5+20x6-S2+A2"\ge" 100

4x1+9x2+11x3+10x4+9x5+10x6-S3+A3=20

and 

x1,x2,x3,x4,x5,x6,A1,A2,A3≥0






Since all Zj-Cj≤0

Hence, optimal solution is arrived with value of variables as :

x1=0.9091,x2=1.8182,x3=0,x4=0,x5=0,x6=0

Min Z=7.2727